Grade 12Magical Mathematics[Interesting Approach]

Find the greatest value of xyz for positive
values of x, y, z subject to the condition
xy + yz + zx = 12.

hjh
10 Years agoGrade 12

2 Answers

hjh
10 Years ago
as the val
jagdish singh singh
10 Years ago
\hspace{-0.74 cm}$ Given $xy+yz+zx=12\;,x,y,z>0.$ Then Using $\bf{A.M\geq G.M}$\\\\ So $\frac{xy+yz+zx}{3}\geq (xyz)^{\frac{2}{3}}\Rightarrow (xyz)\leq 4^{\frac{3}{2}}=8$\\\\and equality hold when $xy=yz=zx.$